MOA-II Obsevation Log

Return to index
Previous day Next day

DATE: 2006/09/08
OBSERVER: S.Nakamura A.Fukui
CONDITION: slightly cloudy -> cloudy -> slightly cloudy -> fine
COMMENT: dark D1399(30sec), D1400(100sec) are OK. CCD-temperature sensor was broken, and now under repair.
Compressor gas pressure: MPa
Expected Observation Time: START 19:23 END 06:06


TARGETRUNTOTAL RUNTOTAL FIELDTOTAL EXPOSURE (sec)
GB4120-4142231344020
LMC613-6131111100
SMC201-20224400
ALERT-000
GRB-000
TOTAL---261495520 (01h32m00s)


GB FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
66652145522455
Field13Field14Field15Field16Field17Field18Field19Field20Field21Field22Field23
44454443350


LMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
110001111111
Field13Field14Field15Field16
1100


SMC FIELD STATISTICS

Field1Field2Field3Field4Field5Field6Field7Field8Field9Field10Field11Field12
111100000000


SPECIAL TARGET NAME

CategoryTargetNrun


CHECK SHEET

TimeStatusWeather
18:30waitingslightly cloudy
19:18observingslightly cloudy, start observin g
19:30waitingstop due to cloudy
22:30observingobserving restart
23:30observingfine
00:30observingfine
01:30observingfine
02:30waitingcloudy, DARK
03:00observingobserving restart, slightly cloudy
03:10waitingstop due to cloudy
04:00waitingcloudy
05:00waitingovercast
06:00waitingovercast


CCD TEMPERATURE Previous day Next day




VACUUM Previous day Next day




TELESCOPE TEMPERATURE Previous day Next day




Return to index

Previous day Next day